OpenStudy (anonymous):

log (1/81) to the base 9 equals to?

OpenStudy (anonymous):

Recall that the log of some function log_a of some number n is some number x, then a^x =n (_ is base) In your case log_9 of (1/81) is 9^x =1/81 recall that an integer l to its negative power w is 1/l^w so for instance 9^-1= 1/9 and 9^-2=1/9^2 or 1/81, and that means that x = -2 and so the answer is -2.